OpenStudy (anonymous):

lim x->o (1-cosx)/x sqrt

myininaya (myininaya):

This seems to be a bit weird sqrt of what?

OpenStudy (anonymous):

it's the new way...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

LOL @myininaya I thought It was \[\LARGE \lim_{x\to 0}{1-\cos x \over \sqrt x}\]

OpenStudy (anonymous):

the lower is x^2

OpenStudy (anonymous):

\[\LARGE \lim_{x\to 0}{1- \cos x\over x^2}\] ?

OpenStudy (anonymous):

yes yes

myininaya (myininaya):

do you l'hospital?

myininaya (myininaya):

i mean know?

OpenStudy (anonymous):

yes. but the question say use limit rule. don't use l'opitall's rule

OpenStudy (anonymous):

i've tried to use factorization. but can't get it

myininaya (myininaya):

ok \[\lim_{x \rightarrow 0}\frac{1-\cos(x)}{x^2} \cdot \frac{1+\cos(x)}{1+\cos(x)}\] \[\lim_{x \rightarrow 0}\frac{1-\cos^2(x)}{x^2(1+\cos(x))}=\lim_{x \rightarrow 0}\frac{\sin^2(x)}{x^2(1+\cos(x))}\] need more help?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(x)}{x} \cdot \lim_{x \rightarrow 0}\frac{\sin(x)}{x} \cdot \lim_{x \rightarrow 0}\frac{1}{1+\cos(x)}\]

OpenStudy (anonymous):

the answer should be?

myininaya (myininaya):

so you got this right ?

OpenStudy (anonymous):

actually no. what technique is this?

myininaya (myininaya):

you wanted to use algebra and limit laws...

myininaya (myininaya):

do you have any questions with the steps i performed?

OpenStudy (anonymous):

is the ans 1/2?

OpenStudy (anonymous):

you're supposed to know this rule: \[ \lim_{x\to0}{\sin x\over x}= ?\] what should be instead of "?"

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so there's nothing to get confused of :) .. @myininaya solved it, you just had to substitute :)

OpenStudy (anonymous):

thank you